OpenStudy (anonymous):

how to find out the value of sec(-960) 'manually'.and how do we decide the sign of the answer at the end?

OpenStudy (anonymous):

first divide by 360 -960/360 = 2 remainder - 240 -240 degrees is equivalent to 120 degrees sec = 1/cos cos 120 degrees = - cos 60 degrees = -1/2 (second quadrant cos is negative) sec = 1 /-1/2 = -2 sec (-960) = -2

OpenStudy (anonymous):

can you explain the general method,that i can apply to all functions of any kind of angle?